Solve the equation 2z^4 -14z^3+33z^2-26z+10 = 0, zeC [where C denotes the set of complex numbers] given that it has root 3+i
1
Expert's answer
2020-11-08T19:08:22-0500

"\displaystyle\\

2z^4 -14z^3+33z^2-26z+10 = 0. \\

\textsf{By the fundamental theorem of}\\
\textsf{algebra, if}\, 3 + i \, \textsf{is a root, its conjugate,}\\
3 - i \, \textsf{is also a root of the polynomial.}\\

(z - (3 + i))(z - (3 - i))\, \textsf{is a factor of the polynomial}\\
(z - (3 + i))(z - (3 - i)) \\= z^2 - (3 + i + 3 - i)z + (3 + i)(3 - i) 
\\=z^2 - 6z + 10 \\



\textsf{By long division}\\

z^2 - 6z + 10 | 2z^2 - 2z + 1\\

2z^4 -14z^3+33z^2-26z+10 \\

-(2z^2 - 12z^3 + 20z^2)\\

-2z^3 + 13z^2 -26z + 10\\

-(-2z^3 + 12z^2 - 20z)\\

z^2 - 6z + 10\\

-(z^2 - 6z + 10)\\

\underline{0}\\


\therefore (z^2 - 6z + 10)(2z^2 - 2z + 1) = 0\\

2z^2 - 2z + 1 = 0\\

\begin{aligned}
z &= \frac{2 \pm \sqrt{4 - 8}}{4} \\&= \frac{2 \pm 2i}{2} = 1 \pm i
\end{aligned}\\



z = 1 + i, 1 - i\\



\therefore z = 1 + i, 1 - i, 3 + i, 3 - i"